# Wireless Network Topology Inference: A Markov Chains Approach

### v1

@@ Abstract
In this work, we address the problem of inferring the topology of a wireless network using limited observational data. Specifically, we assume that we can detect when a node is transmitting, but no further information regarding the transmission is available. We propose a novel network estimation procedure grounded in the following abstract problem: estimating the parameters of a finite discrete-time Markov chain by observing, at each time step, which states are visited by multiple “anonymous” copies of the chain. We develop a consistent estimator that approximates the transition matrix of the chain in the operator norm, with the number of required samples scaling roughly linearly with the size of the state space. Applying this estimation procedure to wireless networks, our numerical experiments demonstrate that the proposed method accurately infers network topology across a wide range of parameters, consistently outperforming transfer entropy, particularly under conditions of high network congestion.
@@ 1 Introduction
Inferring the topology of a network from limited information has numerous applications across various fields, including the discovery of gene interactions in computational biology, the analysis of climate dynamics in environmental science, and the investigation of functional brain structure in neuroscience [net_inf_survey]. Additionally, in defense and security contexts, understanding the structure of an adversary’s network can provide valuable insights for decision-making.
@@
Previous work, particularly in the context of wireless networks [TR13, LC17, TG20, liu2022topology], has focused on detecting links by computing causality measures between time series, such as Granger causality [granger69] and transfer entropy [schreiber00]. While these approaches have demonstrated reasonable accuracy, particularly in simulations, they remain poorly understood. For example, it is unclear in what ways the inferred network approximates the actual network structure.
@@ 1.1 Our Contribution
We address the problem of estimating the topology of a wireless network from limited information. Specifically, we monitor a network of wireless devices, assuming we can detect when a device is transmitting, but not the target of the transmission. Such data could be collected via a network of radio frequency sensors.
@@
Our primary theoretical contribution is the development of a consistent estimator. We prove that by observing the trajectories of the chains for a sufficient duration, we can construct an estimated transition matrix $\hat{P}$ that approximates the true matrix $P$ arbitrarily well in the operator norm. Crucially, this approximation in operator norm provides insights into the global structure of the network, such as the presence of clusters and bottlenecks [STsparse].
We apply this estimation procedure to wireless networks simulated using the ns-3 network simulator [ns3]. Our experimental results show that the proposed method is competitive with state-of-the-art techniques. Notably, we demonstrate that the estimated topology closely approximates the true topology for various networks with differing structures and sizes. Even in cases where the number of observed transmissions is too small for most links to be detected, our procedure successfully recovers structural information about the network, such as the presence of bottlenecks.
@@ 1.2 Related Work
From a theoretical standpoint, our work relates to the broader problem of estimating Markov chain parameters from observed trajectories, an area that has garnered significant attention in recent years [hsu19, Wolf19, WolfKont24]. Our algorithm builds upon techniques developed by Wolfer and Kontorovich [Wolf19]. However, prior work typically focuses on scenarios where only a single Markov chain is observed, which simplifies the estimation task but is less applicable to modelling topology inference in wireless networks. In the single-chain case, observing consecutive transmissions at nodes $u$ and $v$ directly implies the existence of a link between them, a property that does not hold in our model.
There is a significant body of literature on wireless network topology inference (see, e.g.,  [TR13, LC17, salmond19, TG20, mehrotra2021minimax, liu2022topology, du2023network]). However, to the best of our knowledge, our work is the first to provide theoretical guarantees for the estimation procedure, albeit under a simplified model of network dynamics. While this model may not fully capture the complexities of real wireless networks, we believe it encapsulates the key challenges inherent to the inference task.
@@ 2 Problem setup
For technical reasons, we assume $P$ has a unique stationary distribution $\pi$, i.e., $\pi P=\pi$. This is a reasonable assumption in our context since it essentially means the network is (strongly) connected. Furthermore, we assume the Markov chains are stationary, i.e., $X^{(i)}_{0}\sim\pi$ for all $1\leq i\leq k$. This technical assumption is made to simplify our analysis, but is not necessary and could be dropped with additional technical work (without modifying the proposed algorithm).
We assume $P$ and $\pi$ are unknown to us. Indeed, our goal is to estimate $P$. What we are able to observe is, for each time step $t=0,\dots,T$, how many chains are currently visiting a given vertex $v\in V$. In other words, for any $t=1,\dots,T$ and $v\in V$, we can observe the quantity $S_{t,v}=|\{i\colon X^{(i)}_{t}=v\}|$ for any $t\in\mathbb{N}_{+},v\in V$. We are not able, however, to observe if a specific chain is visiting $v$ at time $t$.
@@ 3 Estimation procedure
$N(u)$ =|$=\left|\{i\in[k],0\leq t<T\colon X^{(i)}_{t}=u\}\right|.$
In other words, $N(u,v)$ represents the number of times we have observed $u$’s transmission being immediately followed by $v$’s, while $N(u)$ counts the total number of times $u$ is transmitting. Notice that $N(u)=\sum_{v\in V}N(u,v)$. Moreover, these quantities can be computed having access only to $S_{t,v}$ for $0\leq t\leq T,v\in V$.
@@
The term $(k-1)\widehat{\Pi}$ is a correction necessary because nodes that are visited by the chains more frequently will have larger entries in $M$.
Notice that computing $\hat{P}$ is computationally efficient and can be done in an online manner: we do not have to store the entire time series matrix $TS$, we can just update $N(u,v)$, $N(v)$, and $\hat{P}$ at each time step.
Returning to the wireless network setting, $\hat{P}(u,v)$ can be interpreted as our estimated probability that $v$ will transmit immediately after $u$ has transmitted. While $\hat{P}(u,v)$ is strongly correlated to the likelihood $u$ and $v$ are connected by a link, it is not a measure of how likely a link between $u$ and $v$ exists. To this end, we suggest using an estimator for the Laplacian associated to $P$. For a transition matrix $P$ with stationary distribution $\pi$, the Laplacian $L$ of $P$ is the matrix with entries $L(u,v)=P(u,v)\sqrt{\pi(v)/\pi(u)}$. We recall that any Markov chain can be interpreted as a random walk on a weighted directed graph, where $L(u,v)$ represents the weight on the edge $(u,v)$ [markovmixing]. Therefore, to infer if a link is present or not, we will use the estimated matrix $\hat{L}$, defined as
@@
Finally, when we know that links in the observed network are bidirectional, it is natural to consider a symmetrisation of the Laplacian instead:
@@
We conclude this section by underlining that our estimator only records transmissions happening in two consecutive time intervals. This means it’s important to choose an appropriate interval length, otherwise network delays might severely impact our estimator. In our simulations, however, this constraint has not been detrimental to our results. The situation might be more complicated in real world settings. We plan to study how to bypass these potential issues in future work.
@@ 4 Theoretical guarantees
The theorem essentially tells us that $\hat{P}=M-(k-1)\widehat{\Pi}$ is a good estimator for $P$. In particular, by observing the network for a large enough time, we can approximate $P$ arbitrarily well in the operator norm. This approximation preserves important quantities of the network. For example, it approximately preserves its stationary distribution; this tells us how often, on average, nodes are visited [markovmixing], which corresponds how often a node is transmitting in our model. It also preserves its spectral gap, which tells us how fast a Markov chain with transition matrix $P$ converges to its stationary distribution from a worst case starting point [MonteTetali]. Moreover, it preserves important connectivity properties of the networks, such as its “cut-structure”, that is, the structure of clusters and bottlenecks in the network [PSZ, STsparse].
@@
What this approximation doesn’t necessarily preserve are the individual links in the network. For example, consider a transition matrix of a random walk on a complete graph of $n$ vertices. Now consider an Erdos-Renyi random graph $G(n,p)$: that is, a graph of $n$ vertices where we connect any two vertices with probability $p$. As long as $p\gg log(n)/n$, it is well-known that the transition matrix of a random walk on $G(n,p)$ well approximates (in the operator norm) the transition matrix of the complete graph [STsparse]. However, the $G(n,p)$ graph might have only $O(n\log(n))$ edges compared to the $\Theta(n^{2})$ edges of the complete graph, i.e., only a small fraction of the edges are preserved. Indeed, it is possible to construct examples where Theorem 1 guarantees a good approximation with a number of samples that is much smaller than the number of edges in the graph: certain links wouldn’t have been used at all and, therefore, would not be detectable.
@@
We now discuss the sample complexity of our procedure. The number of samples (or time steps) required depends on several parameters. In particular, it is inversely linear in the spectral gap $\lambda(P)$. The spectral gap is a crucial parameter in the analysis of Markov chains: its reciprocal, $1/\lambda(P)$ provides a bound on the mixing time of $P$, i.e., the time it takes for a Markov chain with transition matrix $P$ to converge to the stationary distribution from a worst case starting point [MonteTetali]. A long as $P$ is irreducible, as in our case, $\lambda(P)>0$. We believe that, in cases where $P$ models a communication network, $1/\lambda(P)$ is relatively small. Furthermore, for natural transition matrices $P$, $1/\pi_{\star}$ is proportional to $n$, the size of the network. A larger value of $1/\pi_{\star}$ would mean, in our model, that there are nodes that transmits very rarely: to approximate $P$, we necessarily need to collect enough information about these nodes, increasing the time required. Finally, $\frac{\pi^{\star}}{\pi_{\star}}$ can be interpreted as the ratio between the largest and the smallest frequencies of transmission of a node: if nodes transmit at roughly the same frequency, this parameter becomes order of 1.
@@
The proof of Theorem 1, deferred to Section 7, is inspired by the techniques of [Wolf19]. In particular, it uses matrix concentration inequalities for Markov chains [matrixMarkovBernstein] and martingales [Tropp].
We end this section by briefly discussing the validity of our Markovian model for capturing our task of estimating the topology of a wireless network. Even though communication networks do not satisfy a Markovian assumption, we hope that, on average, the behaviour of such networks can be approximated well enough by a Markovian model. Besides this consideration, however, for our model to capture the behaviour of wireless networks, an additional assumption must be considered. Since multiple chains can concurrently visit a node at any time, our estimation procedure requires to observe the number of chains visiting a node $u$ at a time $t$. In our modelling, however, a chain visiting a node $u$ at time $t$ corresponds to observing node $u$ transmitting a message during the $t$-th time interval. Therefore, $r$ chains visiting $u$ at time $t$ would correspond to $u$ transmitting $r$ messages during the same time interval. This event, besides being very unlikely, cannot even be detected: we assume we can only observe if a node is transmitting in a certain time interval, not how many messages is transmitting. For this reason, our Markovian model can only be realistic if the chance of two or more chains visiting the same node at the same time is very small. Formally, the probability that at a time $t$ two stationary chains visit the same node can be upper bounded by
@@ 5 Experimental results
We have simulated various wireless network topologies with the network simulator ns-3 [ns3]. Each wireless network is constructed by generating $n$ points on the plane. Each point corresponds to a wireless device and two devices are able to communicate with one another if they are at most 50m apart. Different point sets will induce a different network topology. Propagation speed is constant (the speed of light) and path loss is dependent only on the transmission range: if the receiver is within range, it will receive the transmission at 16dbm, else it will not receive a transmission. We use the IEEE 802.11b protocol with a data rate of 1Mbps. We use the OLSR routing protocol to discover and disseminate edge state information. The Hello and TC intervals are 2 and 5 seconds, respectively. We do not send any packets within the first 30 seconds of the simulation to allow time for devices to calculate the initial shortest path between other nodes in the network. We use UDP as the transport layer protocol and each node is implemented as a UDP client. This choice has been made because UDP does not use acknowledgment messages (ACKs), possibly making link detection more difficult. However, we have also performed some simulations with TCP obtaining very similar results.
@@
Our experiments assume the network is bidirectional, and therefore, we will use our estimate $\hat{L}_{\text{sym}}$ for the symmetrised Laplacian. We will compare our estimate with one derived from transfer entropy. More precisely, we will consider the matrix $M_{\text{TE}}$ where the $(i,j)$-entry is equal to the transfer entropy from the time series associated with $i$ to the time series associated with $j$. To make the comparison even, we will further symmetrised this matrix by taking the average between this matrix and its transpose.

### v2

@@ Abstract
We address the problem of inferring the topology of a wireless network using limited observational data. Specifically, we assume that we can detect when a node is transmitting, but no further information regarding the transmission is available. We propose a novel network estimation procedure grounded in the following abstract problem: estimating the parameters of a finite discrete-time Markov chain by observing, at each time step, which states are visited by multiple “anonymous” copies of the chain. We develop a consistent estimator that approximates the transition matrix of the chain in the operator norm, with the number of required samples scaling roughly linearly with the size of the state space. Applying this estimation procedure to wireless networks, our numerical experiments demonstrate that the proposed method accurately infers network topology across a wide range of parameters, consistently outperforming transfer entropy, particularly under conditions of high network congestion.
{IEEEkeywords}
Markov chains, topology inference, wireless networks.
@@ 1 INTRODUCTION
Inferring the topology of a network from limited information has numerous applications across various fields, including the discovery of gene interactions in computational biology, the analysis of climate dynamics in environmental science, and the investigation of functional brain structure in neuroscience [1, 2, 3]. Additionally, in defense and security contexts, understanding the structure of an adversary’s network can provide valuable insights for decision-making.
@@
Previous work, particularly in the context of wireless networks [4, 5, 6, 7], has focused on detecting links by computing causality measures between time series, such as Granger causality [8] and transfer entropy [9]. While these approaches have demonstrated reasonable accuracy, particularly in simulations, they remain poorly understood. For example, it is unclear in what ways the inferred network approximates the actual network structure.
@@ 1.1 Our Contribution
We address the problem of estimating the topology of a wireless network from limited information. Specifically, we monitor a network of wireless devices, assuming we can detect when a device is transmitting, but not the target of the transmission. Such data could be collected via a network of radio frequency sensors (see Fig. 1 for a schematic depiction).
Figure 1: A network of wireless devices (blue circles) monitored by a collection of radio frequency sensors (green squares). Whenever a wireless device is transmitting, nearby sensors will measure an associated power bursts. Our goal is to infer the topology of the network using only the sensors’ measurements.
@@
Our primary theoretical contribution is the development of a consistent estimator. We prove that, by observing the trajectories of the chains for a sufficient duration, we can construct an estimated transition matrix $\hat{P}$ that approximates the true matrix $P$ arbitrarily well in the operator norm. Crucially, this approximation in operator norm provides insights into the global structure of the network, such as the presence of clusters and bottlenecks [10].
We apply this estimation procedure to wireless networks simulated using the ns-3 network simulator [11]. Our experimental results show that the proposed method is competitive with state-of-the-art techniques. Notably, we demonstrate that the estimated topology closely approximates the true topology for various networks with differing structures and sizes. Even in cases where the number of observed transmissions is too small for most links to be detected, our procedure successfully recovers structural information about the network, such as the presence of bottlenecks.
@@ 1.2 Related Work
From a theoretical standpoint, our work relates to the broader problem of estimating Markov chain parameters from observed trajectories, an area that has garnered significant attention in recent years [12, 13, 14]. Our algorithm builds upon techniques developed by Wolfer and Kontorovich [13]. Prior work, however, typically focuses on scenarios where only a single Markov chain is observed, which simplifies the estimation task but is less applicable to modelling topology inference in wireless networks. In the single-chain case, observing consecutive transmissions at nodes $u$ and $v$ directly implies the existence of a link between them, a property that does not hold in our model.
There is a significant body of literature on wireless network topology inference (see, e.g.,  [4, 5, 15, 6, 16, 7, 17]). However, to the best of our knowledge, our work is the first to provide theoretical guarantees for the estimation procedure, at least under a simplified model of network dynamics. While this model may not fully capture the complexities of real wireless networks, we believe it encapsulates the key challenges inherent to the inference task.
@@
1.3 Organisation
We introduce the setup and the formal model underpinning our inference procedure in Section 2, which then is described in Section 3. In Section 4 we state and discuss our main theoretical result, while its proof is deferred to Section 7. We showcase and discuss our experimental results in Section 5. In particular, we first consider an idealised scenario in which we know exactly when a given node is transmitting (Section 5 5.1) and we later discuss how to acquire this information using radio frequency sensors (Section 5 5.2). We end with a discussion of possible future work in Section 6. Code and datasets to reproduce our experiments are available at [18].
@@ 2 PROBLEM SETUP
We remark that, in practical scenarios, we do not usually have access to perfect information about when a device is transmitting. Rather, we can only obtain an approximation of the time series matrix, for example, by deploying a set of radio frequency sensors in the environment. We expand on how RF sensors can be used to approximate this time series matrix in Section 5 5.2.
@@
For technical reasons, we assume $P$ has a unique stationary distribution $\pi$, i.e., $\pi P=\pi$. This is a reasonable assumption in our context since it essentially means the network is (strongly) connected, i.e., any two nodes of the network can communicate with each other (without necessarily being connected by a link). Furthermore, we assume the Markov chains are stationary, i.e., $X^{(i)}_{1}\sim\pi$ for all $1\leq i\leq k$. This technical assumption is made to simplify our analysis, but is not necessary and could be dropped with additional technical work (without modifying the proposed algorithm).
We assume $P$ and $\pi$ are unknown to us. Indeed, our goal is to estimate $P$. What we are able to observe is, for each time step $t=1,\dots,T$, how many chains are currently visiting a given vertex $v\in V$. In other words, for any $t=1,\dots,T$ and $v\in V$, we can observe the quantity $S_{t,v}=|\{i\colon X^{(i)}_{t}=v\}|$. We are not able, however, to observe if a specific chain is visiting $v$ at time $t$.
@@ 3 ESTIMATION PROCEDURE
$N(u)$ =|$=\left|\{i\in[k],1\leq t<T\colon X^{(i)}_{t}=u\}\right|.$
In other words, $N(u,v)$ represents the number of times we have observed $u$’s transmission being immediately followed by $v$’s, while $N(u)$ counts the total number of times $u$ is transmitting. Notice that $N(u)=\sum_{v\in V}N(u,v)$. Moreover, these quantities can be computed having access only to $S_{t,v}$ for $1\leq t\leq T,v\in V$.
@@
The term $(k-1)\widehat{\Pi}$ is a correction necessary because nodes that are visited by the chains more frequently will have larger entries in $M$; in particular, if the chains often visit two nodes $u$ and $v$, the corresponding entry $M(u,v)$ will be large even though there might not be an edge between $u$ and $v$. Indeed, we expect that, on average, $M(u,v)\approx(k-1)\widehat{\Pi}$ whenever $u$ is not connected to $v$.
Computing $\hat{P}$ is computationally efficient and can be done in an online manner: we do not have to store the entire time series matrix $TS$, we can just update $N(u,v)$, $N(v)$, and $\hat{P}$ at each time step. This can be done in time $O(n^{2}T)$ in total.
Returning to the wireless network setting, $\hat{P}(u,v)$ can be interpreted as our estimated probability that $v$ will transmit immediately after $u$ has transmitted. While $\hat{P}(u,v)$ is strongly correlated to the likelihood $u$ and $v$ are connected by a link, it is not a measure of how likely a link between $u$ and $v$ exists. In particular, if a node $u$ has many outgoing links, entries in the corresponding row $\hat{P}(u,\cdot)$ will necessarily be small.
To overcome this issue, we suggest using an estimator for the Laplacian associated to $P$. For a transition matrix $P$ with stationary distribution $\pi$, the Laplacian $L$ of $P$ is the matrix with entries $L(u,v)=P(u,v)\sqrt{\pi(v)/\pi(u)}$. We recall that any Markov chain can be interpreted as a random walk on a weighted directed graph, where $L(u,v)$ represents the weight on the edge $(u,v)$ [19]. Therefore, to infer if a link is present or not, we will use the estimated matrix $\hat{L}$, defined as
@@
When we know that links in the observed network are bidirectional, it is natural to consider a symmetrisation of the Laplacian [20] instead:
@@
We conclude this section by underlining two issues that arise when applying our methodology to infer the topology of real-world wireless networks. First of all, notice that our estimator only records transmissions happening in two consecutive time intervals. This means it is important to choose an appropriate interval length, otherwise network delays might severely impact our estimator. In our simulations, however, this constraint has not been detrimental to our results. The situation might be more complicated in real world settings. Secondly, our estimator depends on the number $k$ of Markov chains that are concurrently active. In practice, $k$ can be interpreted as the average number of devices that transmit in the same time interval. We will estimate this parameter as the ratio between the number of nonzero entries in the time series matrix $TS$ and the number of time-intervals $t$ such that there has been a transmission both at interval $t$ and at the following interval $t+1$ (this is because our estimator $\hat{P}$ only considers consecutive transmissions).
@@ 4 THEORETICAL GUARANTEES
The theorem essentially tells us that $\hat{P}=M-(k-1)\widehat{\Pi}$ is a good estimator for $P$. In particular, by observing the network for a large enough time, we can approximate $P$ arbitrarily well in the operator norm. This approximation preserves important quantities of the network. For example, it approximately preserves its stationary distribution; this tells us how often, on average, nodes are visited [19], which corresponds how often a node is transmitting in our model. It also preserves its spectral gap, which tells us how fast a Markov chain with transition matrix $P$ converges to its stationary distribution from a worst case starting point [20]. Moreover, it preserves important connectivity properties of the networks, such as its “cut-structure”, that is, the structure of clusters and bottlenecks in the network [21, 10].
@@
What this approximation doesn’t necessarily preserve are the individual links in the network. For example, consider a transition matrix of a random walk on a complete graph of $n$ vertices. Now consider an Erdős-Rényi random graph $G(n,p)$: that is, a graph of $n$ vertices where we connect any two vertices with probability $p$. As long as $p\gg log(n)/n$, it is well-known that the transition matrix of a random walk on $G(n,p)$ well approximates (in the operator norm) the transition matrix of the complete graph [10]. However, the $G(n,p)$ graph might have only $O(n\log(n))$ edges compared to the $\Theta(n^{2})$ edges of the complete graph, i.e., only a small fraction of the edges are preserved. Indeed, it is possible to construct examples where Theorem 1 guarantees a good approximation with a number of samples that is much smaller than the number of edges in the graph: certain links wouldn’t have been used at all and, therefore, would not be detectable.
@@
We now discuss the sample complexity of our procedure. The number of samples (or time steps) required depends on several parameters. In particular, it is inversely linear in the spectral gap $\lambda(P)$. The spectral gap is a crucial parameter in the analysis of Markov chains: its reciprocal, $1/\lambda(P)$ provides a bound on the mixing time of $P$, i.e., the time it takes for a Markov chain with transition matrix $P$ to converge to the stationary distribution from a worst case starting point [20]. As long as $P$ is irreducible, as in our case, $\lambda(P)>0$. We believe that, in cases where $P$ models a communication network, $1/\lambda(P)$ is relatively small. Furthermore, for natural transition matrices $P$, $1/\pi_{\star}$ is proportional to $n$, the size of the network. A larger value of $1/\pi_{\star}$ would mean, in our model, that there are nodes that transmit very rarely: to approximate $P$, we necessarily need to collect enough information about these nodes, increasing the time required. Finally, $\frac{\pi^{\star}}{\pi_{\star}}$ can be interpreted as the ratio between the largest and the smallest frequencies of transmission of a node: if nodes transmit at roughly the same frequency, this parameter becomes order of 1.
@@
The proof of Theorem 1, deferred to Section 7, is inspired by the techniques of [13]. In particular, it uses matrix concentration inequalities for Markov chains [22] and martingales [23].
We end this section by briefly discussing the validity of our Markovian model for capturing our task of estimating the topology of a wireless network. Even though communication networks do not satisfy a Markovian assumption, we hope that, on average, the behaviour of such networks can be approximated well enough by a Markovian model. Besides this consideration, however, for our model to capture the behaviour of wireless networks, an additional assumption must be considered. Since multiple chains can concurrently visit a node at any time, our estimation procedure requires to observe the number of chains visiting a node $u$ at a time $t$. In our modelling, however, a chain visiting a node $u$ at time $t$ corresponds to observing node $u$ transmitting a message during the $t$-th time interval. Therefore, $r$ chains visiting $u$ at time $t$ would correspond to $u$ transmitting $r$ messages during the same time interval. This event, besides being very unlikely, cannot even be detected: we assume we can only observe if a node is transmitting in a certain time interval, not how many messages is transmitting. For this reason, our Markovian model can only be realistic if the chance of two or more chains visiting the same node at the same time is very small. Formally, the probability that at some time $t$ two stationary chains visit the same node can be upper bounded by
@@ 5 EXPERIMENTAL RESULTS
We have simulated various wireless network topologies with the network simulator ns-3 [11]. Each wireless network is constructed by generating $n$ points on the plane. Each point corresponds to a wireless device and two devices are able to communicate with one another if they are at most 50m apart. Different point sets will induce a different network topology. We use the IEEE 802.11b protocol with a data rate of 1Mbps. We use the OLSR routing protocol to discover and disseminate edge state information. The Hello and TC intervals are 2 and 5 seconds, respectively. We do not send any packets within the first 30 seconds of the simulation to allow time for devices to calculate the initial shortest path between other nodes in the network. We use UDP as the transport layer protocol and each node is implemented as a UDP client. This choice has been made because UDP does not use acknowledgment messages (ACKs), possibly making link detection more difficult. However, we have also performed some simulations with TCP obtaining very similar results.
@@
Our experiments assume the network is bidirectional, and therefore, we will use our estimate $\hat{L}_{\text{sym}}$ for the symmetrised Laplacian. We will compare our estimate with one derived from transfer entropy. More precisely, we will consider the matrix $M_{\text{TE}}$ where the $(i,j)$-entry is equal to the transfer entropy from the time series associated with $i$ to the time series associated with $j$. To make the comparison even, we will also symmetrise this matrix by taking the average between this matrix and its transpose. Transfer entropy is estimated using the Python library PyInform [24] with embedding dimension $d=5$. We did not optimise such number but found a similar performance for smaller values of $d$, while larger values of $d$ might produce worse results.
5.1 An idealised setting
We first consider an idealised setting that follows the problem setup described in Section 2: we assume we have direct access to the time series matrix $TS$. While we do not expect to have perfect information about the time series matrix $TS$ in practice, we believe useful to know the strengths and limitations of our approach in this idealised scenario.
We fix a constant propagation speed (the speed of light) and path loss is dependent only on the transmission range: if the receiver is within range, it will receive the transmission at 16dbm, else it will not receive a transmission.